OpenStudy (31356):

Through: (0,5), perpendicular to y= 3/5x-1 Please help, medal will be earned.

OpenStudy (31356):

The answer will be in slope intercept form.

OpenStudy (hoblos):

the slope of the perpendicular line is the negative reciprocal of the slope of the first line this means , m * m' = -1 can you calculate the slope of the perpendicular line now?

OpenStudy (31356):

Can you please show me step by step how to do this because I can't find the mistake where I messed up at.Thanks. Also yes, I know that m*m =-1 and the reciprocal.

OpenStudy (hoblos):

if m is the slope of the perpendicular m * 3/5 = -1 m = -1 * 5/3 so m=-5/3 so the equation of the line will be y= -5/3 x +b now you have to replace the point (0,5) in the equation to calculate b

OpenStudy (31356):

Oh right thanks, I figured where I messed up now thank you for your time & explanation!

OpenStudy (hoblos):

any time :)